An asset used in a four-year project falls in the five-year MACRS class (MACRS Table) for tax purposes. The asset has an acquisition cost of $6.8 million and will be sold for $1.46 million at the end of the project. If the tax rate is 23 percent, what is the aftertax salvage value of the asset?
Answer #1

book value of asset = cost - accumulated depreciation

= 6800000 - 6800000 * (0.2+0.32+0.1920+0.1152)

= 1175040

tax on sale = (1460000 - 1175040) * 0.23

= 65540.8

after tax salvage value of asset = 1460000 - 65540.8

= 1394459.2